The WireOne Company manufactures high-quality coated electrical wire in two departments, Weaving and Coating. Materials are introduced at various points during work in the Weaving Department. After the weaving is completed, the materials are transferred into the Coating Department, where specialty plastic coating is applied.

Selected data relating to the Weaving Department during May are given below:
  Production data:
    Kilograms in process, May 1(materials 100%
          complete; conversion 80% complete)
127,500
    Kilograms started into production during May 448,000
    Kilograms completed and transferred to Coating ?
    Kilograms in process, May 31(materials 65%           complete; conversion 30% complete) 68,000
  Cost data:
    Work in process inventory, May 1:
        Materials cost $ 113,475
        Conversion cost $ 132,600
    Cost added during May:
        Materials cost $ 824,415
        Conversion cost $ 437,532
The company uses the weighted-average method.
Required:
1. Compute the equivalent units of production.
Answer #1
WEIGHTED AVERAGE METHOD:
Physical flow schedule:
Units to account for:
Units in BWIP 127500
Units started 448000
Total units to account for 575500
Units accounted for:
Units completed and transferred 507500
Units in EWIP 68000
Total units accounted for 575500
Schedule of equivalent units: Materials Conversion Total
Units completed 507500 507500
Units in EWIP:
=68000*65% = 44200 0
=68000*30% = 0 20400
Equivalent units of production 551700 527900
Cost per equivalent unit:
Cost in BWIP $           1,13,475 $      1,32,600 $       2,46,075
Costs added during the month $           8,24,415 $      4,37,532 $     12,61,947
Total costs for equivalent units $           9,37,890 $      5,70,132 $     15,08,022
Equivalent units of output 551700 527900
Cost per equivalent unit $                  1.70 $              1.08 $               2.78
Cost of goods transferred out $           8,62,750 $   5,48,100 $ 14,10,850
Cost of EWIP $              75,140 $          22,032 $           97,172
Cost reconciliation:
Costs to account for:
BWIP $        2,46,075
Incurred during the period $      12,61,947
Total costs to account for $    15,08,022
Costs accounted for:
Goods transferred out $    14,10,850
EWIP $           97,172
Total costs accounted for $    15,08,022